What is the differnce between heat exchange in a process & the Enthalpy Change??


Why is the Enthalpy Change is an isothermal reversible process 0 even when some work might be done for isothermal expansion of the gas?

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

H=U+PV,ΔH=ΔU+ΔPV=ΔQ-ΔW+nR*ΔT,so change in enthalpy=heat inflow+work done+increase in internal energy

in isothermal reversible process,ΔT=0,implying that ΔPV=0,

therefore,ΔH=ΔU=ΔQ-ΔW

Now, ΔQ=ΔW, therefore ΔH=ΔU=0

Suman Ghosh
12 years ago

for an isothermal process how is ΔPV=0?

i mean work can be done in an isothermal process owing to expansion in volume... for instance, suppose boiling of water into steam... then??